OpenStudy (anonymous):

Factor: x2 - 24x + 144

OpenStudy (anonymous):

x^2 - 24x + 144

mathslover (mathslover):

\[\huge{x^2-24x+144}\] \[\huge{x^2-12x-12x+144}\] \[\huge{x(x-12)-12(x-12)}\] \[\huge{(x-12)(x-12)=(x-12)^2}\]

OpenStudy (callisto):

x^2 - 24x + 144 = x^2 - 2(12)(x) + 12^2 Apply the identity \(a^2 - 2ab + b^2 = (a-b)^2\) where in this case, a=x and b=12 = (x-12)^2

OpenStudy (anonymous):

A. (x - 12)(x + 12) B. (x - 6)(x + 24) C. (x - 12)2 D. (x - 72)2

OpenStudy (anonymous):

so c